L Band Wide Gregorian Dome Receiver Log

Go to Most Recent Entry

Name Email Date Comment
ANS anolla@naic.edu 5/23/2002 IN LAB: The receiver Dewar was cooled for test purpose . Its vacuum was 3.5 un's after 18 hrs of pumping. We started its cool down at the 09:05 A.M. and after an hour (10:00 A.M.) its vac was .1 un / VPG 6 / 16 stage =17.1 / 70 stage =136.2 / OMT= 282
ANS anolla@naic.edu 5/28/2002 IN LAB: System still been tested.Normal readings VPG 2 / 16= 12.9 / 70 55.7 /Return pressure 90. Cryocomp return 95
lswlwray@naic.eduMonday, February 17, 2003 10:00:00INSTALLATION DAY. Brought up receiver to dome, installed in floor. Temoporary bias box mount made
lswlwray@naic.eduTuesday, February 18, 2003Horn support added by Vives
lswlwray@naic.eduWednesday Feb 19 2003Permanent bias box mount made
ansanolla@naic edu.Thursday, February 20, 2003 14:54:32Add circulators to the filter bank.
lswlwray@naic.eduThursday, March 6, 2003 17:21:20Phil taking Tsys and Tcal measurements using correlator. Saw small ripples in Ch A, found problem to be from filter bank.
ansanolla@naic.eduThursday, March 13, 2003 14:34:04Filter bank was modified in order to avoid saturation from SJ aiport radar,
ansanolla@naic.eduTuesday, April 8, 2003 12:48:23System was hooked up for the SETI people.The splitter was added in order to inject noise into the system.
ansanolla@naic.eduTuesday, April 29, 2003 13:07:18The rf cables from the Dewar were replaced with semi rigid cable.Also the plastic power connector was rf shielded and the horn skirt was installed.
ansanolla@naic.eduWednesday, May 7, 2003 14:49:15There it was a rearrengment with the filter bank configuration in order to get rid of the RFI caused in the band due to the different radars operating on it.To see this new configuration go to Lisa's Home Page.
ansanolla@naic.eduTuesday, May 13, 2003 09:12:18The rf amplifiers of the filter bank output and the ones in the polarization box were swap in order to lower down the gain to avoid saturation coming from the different radars rfi. Now we have the Cougars at the filter bank and the Miteqs at the Polarization box.
ansanolla@naic.eduFriday, July 11, 2003 13:53:55The horn test performed on system , existing horn window broken ,fixed.
jrvjrosa@naic.eduTuesday, August 26, 2003 11:25:29The remote temperature reading board in the 70 stage is causing an oscillation in bias card on chann B stage 2.
ansanolla@naic.eduFriday, August 29, 2003 11:27:41The remote cryo reading board got burned while looking for an unstability problem in the second bias stage of channel B.So board was removed from ckt.No cryo temp reading from these recvr. until further notice.
ansanolla@naic.eduWednesday, September 10, 2003 12:42:17The remote cryo temp reading is back , but there is still problems with the system. The Lakeshore meter apperently is feeding back a -13 vdc to the analog sw on the cryo temp board.The system was left without the Lakeshore meter for a while in order to be able to have the remote cryo temp reading.
ansanolla@naic.eduTuesday, September 23, 2003 08:43:26The post amp chassis removed in order to work with a problem with the Lakashore meter , JRV De Roman found that there was a short between and rf cable and the meter sw at the front of the chassis.
ansanolla@naic.eduWednesday, October 8, 2003 12:49:50The remote cryo reading was fix. The sw board was bad .It was replaced with an spare one . So remote cryo reading back to normal.
ansanolla@naic.eduWednesday, October 15, 2003 12:31:51Channel A buffer installed for ALFA recvr test , using fiber optic lines to send down rf (FOL #4 been used).
ansanolla@naic.eduMonday, October 27, 2003 13:21:21Shutter removed from horn after dual beam run.
ansanolla@naic.eduTuesday, March 16, 2004 13:10:18The cryo lines were replaced. They were old lines ( 1/2") 8(R) & 9(S). Now are new lines (3/4") 15(R)& 16(S) .
lswlwray@naic.eduThursday, March 25, 2004 15:21:56Filterbank: found a cable in the 1320 MHz waveguide HP filter connection that switched Ch A with B. Remedied.
ansanolla@naic.eduThursday, May 6, 2004 08:18:26April 19 - 20 Maintenance working in dome to get ready for ALFA April 21: ALFA is installed April 23 or so: used LBW to test ALFA.
ansanolla@naic.eduMonday, May 17, 2004 11:35:26System down in order to check the cal diference.
cheojrosa@naic.eduThursday, May 20, 2004 10:38:091-bias box replaced the switches with safety switches on the power, heater and rf amps in the bias box. added a 0.1uf cap to the mux board constant current source to kill a 123khz oscillation, and a diode in series with the +15v line at the input to the board to isolate the 15v supply from the TTL signal input when the power is off. added bypass cap C504 through C804 (0.1uf in parallel with 1uf) to the bias board ( these units are on the schematic, but not on the board). added hardware to prevent the power cord from being accidentally pulled off from the chassis. 2- postamp replaced the power switch with a safety switch on the post amp. added hardware to prevent the power cord from being accidentally pulled off from the chassis.
lswlwray@naic.eduTuesday, June 15, 2004 08:23:28June 14: Took down postamp chassis to look for high Tsys problem.
lswlwray@naic.eduMonday, June 21, 2004 08:36:53Friday June 18: Replaced postamp chassis.
lswlwray@naic.eduTuesday, June 22, 2004 13:58:06Troubleshooting downstairs. Compared sqare-law detected noise on both Ch A and B and they are comparable. Not the 20% difference we expect if we believe the Tsys values. Problem points to the cal itself.
lswlwray@naic.eduWednesday, June 23, 2004 16:24:18Testing on the Tsys problem. Exchanged the cal cables (including the phase shifters on both Cal A and B) with some lab cables (orange with black ends, 3' LMR 240 types). Re-did the Tsys, and the problem still occurs (i.e. Tsys is 28K/32K for A/B).
lswlwray@naic.eduThursday, June 24, 2004 15:52:36Testing with a +13 dBm 2380 signal into a waveguide near the transmitter. The signal travels down the waveguide and out the SB Tx horn. Testing the LBW and ALFA to see how much of this signal gets back into the LNAs. Tested the mesh cover for LBW today. Results to be posted soon...
ansanolla@naic.eduThursday, July 1, 2004 10:40:45Cal chassis down , been checked for chan B cal fluctuations.
lswlwray@naic.eduThursday, July 1, 2004 15:07:28Postamp/cal box taken down for the day to test out the calB section, as the Tsys for Ch B is 5K higher than A and has been moving for the last few months. Replaced the cal cables with new ones, removed the SETI couplers, checked things out. Cleaned some dirty connections.
ansanolla@naic.eduFriday, July 2, 2004 11:21:34System down due to power shut down (shorted power line at the Rotary Joint).
ansanolla@naic.eduSaturday, July 3, 2004 16:25:31System back to normal , after cool down.
ansanolla@naic.eduMonday, October 11, 2004 15:54:37System was down due to power failure on Saturday , was cooled down by BG and Ganesh.
ansanolla@naic.eduTuesday, October 19, 2004 15:51:23Filter 3B of filter bank was not switching , it was due that the control signal coming from the driver had a broken cable.
ansanolla@naic.eduWednesday, November 3, 2004 14:24:28There is a cryo leak at the LBW cryo system , it is more than 10 PSI/day . Vives is checking it.
ansanolla@naic.eduThursday, November 4, 2004 11:44:02J. Vives found the return line at the cable wrap leaking. Replaced with an spare one.System is ok no leak.
ansanolla@naic.eduTuesday, November 9, 2004 10:56:43System warmed up in order to replace its cross head on 11/12/2004.
ansanolla@naic.eduFriday, November 12, 2004 12:54:17Coco replaced the crosshead , it was ED 030 now is ED 027. System been pumped , cryo system will be turned on tomorrow.
ansanolla@naic.eduMonday, November 15, 2004 13:39:12System back to normal after replacing dewar cryohead.Tsys is chan A 31.5k and B 29.9k.
ansanolla@naic.eduTuesday, November 30, 2004 13:12:23LBW filter bank channel B was not working properly , it was due that the rf cables were not well tight in the 1320 Mhz HP filter. Also we found that the rf cables were in bad shape , so we replaced all of them in both channels on this HP filters.
ansanolla@naic.eduTuesday, November 30, 2004 13:18:40The calibration system was not working properly in all the receivers at the Dome . It was due that the post amp chassis for the CBH and XB were off. These was loading the other recvrs calibrators.
ansanolla@naic.eduMonday, December 6, 2004 14:46:24Post & bias chassis down for maint.
ansanolla@naic.eduFriday, December 10, 2004 15:54:49System back to normal. Tsys was channel A=26k & B=28k.
ansanolla@naic.eduWednesday, January 12, 2005 13:07:48Cryocomp ED 013 was swap with cryocomp ED 016. It was done by JVR in order to clean out the mess at the Dome cryocomp area.
ansanolla@naic.eduThursday, January 13, 2005 12:53:27There is a He leak at the cryo system ,when we swap cryocomp ED 013 with ED 016 , we left it charged to 280.
ansanolla@naic.eduFriday, January 14, 2005 11:44:35Ther is a leak on system , was 280 yesterday and today was 260. I look for it and only found a medium leak at the inline filter at the crosshead.
ansanolla@naic.eduTuesday, February 1, 2005 11:39:53Noise injection board replaced.
ansanolla@naic.eduWednesday, March 2, 2005 11:11:58Channel B 1370 Mhz filter had a misstreaded n/sma connector at bottom of it,was replaced. Also channel A pos 6 filter bank had a loose sma connector inside , it was the one between filter and ext connector.Forgot to log in these information last week.
ansanolla@naic.eduTuesday, May 17, 2005 12:50:28There it is an unstability on channel B.It was pinpointed that it comes from inside the Dewar.It ran away when you turn off the crosshead.So BG banged the Dewar itself an walla , something is loose inside the Dewar.So due to this condition Ganesh is requesting for time in order to fix it.
jrosajrosa@naic.eduFriday, June 3, 2005 14:48:33on 31 may 05 (tuesday)opened the lband receiver. found chan b cal cable and chan a output rf cable with bandly soldered connectors on the 300 kelvin end. the two rf input cables were over tightened by the 16 degree strap. the chan a rf output and the chan b cal cables were repaired. two new rf input cables were made. on 2 june 05 the cables were installed and the vacuum pump connected. on 3 june 03 cool down was started.
ansanolla@naic.eduWednesday, July 13, 2005 16:26:33There were broken control wires at the filter bank.So the 1370 Mhz channel B waveguide filter was not going due to this.
ansanolla@naic.eduFriday, July 22, 2005 12:51:37Tha micronetics NS-213 cal diode of channel A died. It was replaced with an spare one whose sn # is 7002.System needs to recalibrate channel A cals.
W. Genterwgenter@naic.eduFriday, July 22, 2005 15:33:21Noise diode A failed yesterday afternoon. At this time all cal's are h90cal or l90cal.
W. Genterwgenter@naic.eduThursday, July 28, 2005 10:49:28Reported instability in Ch B last evening. Found loose connector in CH B 1290 HPF.
ansanolla@naic.eduFriday, August 26, 2005 13:02:00Due to power failure on 8/25/2005 system warmed up. Been cooled.
ansanolla@naic.eduTuesday, August 30, 2005 16:01:14System back to normal.
ansanolla@naic.eduThursday, October 20, 2005 13:11:59A new safety divice added to the crosshead driver power sw in order to avoid to turn it off accidentally.
ansanolla@naic.eduTuesday, November 1, 2005 14:49:56The filter bank was checked in order to find an intermittence on chnn B 3. Nothing was found loose in that channel.In an overall check out , filters 3 & 4 on channel A had loose connectors , all others were ok.
ansanolla@naic.eduWednesday, November 16, 2005 10:59:26System down ,post amp chassis removed in order to install new cal diodes.
ansanolla@naic.eduWednesday, November 23, 2005 10:42:06System back to normal yesterday (11/22/2005). It was down due to the installation of new cal diodes on it (noise com ones).Was calibrated today (11/23/2005).
ansanolla@naic.eduWednesday, December 21, 2005 11:40:37Tsys was checked , due that in yesterday (12/20/2005 / AST 18:00)Tsys check out the data shown that the receiver cal was bad ,crazy negative #.Today we performed several Tsys check out and all were OK.Program failure as it was at the 430Mhz Dome receiver in the past? An eye on it.
ansanolla@naic.eduMonday, March 6, 2006 13:24:52The cal problem of channel B still there . Try this morning to pinpoint its cause but after several times trying to locate it, we didn't.Every test pinpoint in our judge that it comes inside the Dewar
ansanolla@naic.eduThursday, June 15, 2006 13:23:59System shut down for maintenance since 6/14/1006 AST 04:00.Replacing cryo comp ED 016 by cryo comp ED 007. Also Dewar was open to check for an unstability on channel B.Loose sma connector found in the cal of channel B , it was inside the Dewar feed thru connector of cal B.Been cooled presently.
ansanolla@naic.eduFriday, June 16, 2006 11:08:31System back into service since these morning.Also we found yesterday in the afternoon than the Dewar vacuum was left reading 1.5 uns before noon . The problem was pinpointed to the big vacuum pump , it was replaced by one of the small ones and vacuum returned to 1.5 uns.Big pump been cleaned by running it.
ansanolla@naic.eduFriday, June 16, 2006 11:10:52Forgot to mention in the above date that vaccum was fond reading 70 uns at the Dewar in the afternoon.
ansanolla@naic.eduThursday, July 6, 2006 12:54:19System was calibrated today by Felipe.
ansanolla@naic.eduFriday, September 1, 2006 11:31:44LBW cryo system hose 15 was removed in order to replace an ALFA 2 cryo system leaky hose at the cable wrap.So receiver warmed up for around 1 hour,pump on Dewar for recooling.
ansanolla@naic.eduTuesday, November 21, 2006 14:38:53Cross head noisy.
ansanolla@naic.eduWednesday, January 17, 2007 13:14:45Cryo compressor (ED 007) was cycling due to bad cooling fan , cryo temps were up due to this. So we took the CBH receiver cryo comp (ED 010)to run this system.
answhoknows@naic.eduThursday, February 1, 2007 14:40:14Cryo comp ED 016 installed to run the LBW.ED 010 back to CBH.
ansanolla@naic.eduTuesday, February 6, 2007 11:35:52The cryohead got stocked.So warming up system in order to replace the cryohead.It will take two days at least to got the system back to normal.
ansanolla@naic.eduWednesday, February 7, 2007 16:04:23System being cooled down. Cryo head replaced.Also contamination test of cryo comp ED 016 performed by JVR , it came with negative results.
ansanolla@naic.eduThursday, February 8, 2007 15:47:56System back to normal.New crosshead on it (ED 037 ).
ansanolla@naic.eduTuesday, February 20, 2007 14:59:57Channel 6 A&B patch panel cables at filter bank replaced. Are a match pair.
ansanolla@naic.eduTuesday, March 20, 2007 12:37:37System down until further notice for maintenance.
ansanolla@naic.eduThursday, March 29, 2007 12:40:10Carlos found that a cryo head rocker was loose. He is rechecking the cryo head. Is the last one that was refurbish by Rudy crew.
ansanolla@naic.eduFriday, March 30, 2007 16:01:49New cal & rf cables were made for the system (Dewar ones).
ansanolla@naic.eduMonday, April 23, 2007 12:59:51System was calibrated this morning.
CRVcrios@naic.eduWednesday, January 2, 2008 14:27:52Vacuum pump connected to dewar and pumping. Dic.5th,2007
CRVThursday, January 3, 2008 09:28:51Dewar cooldown Started on Dicember 8th,2007
CRVThursday, January 3, 2008 09:29:18Dewar cold on Dicember 9th,2007
crvThursday, January 3, 2008 10:39:19Having problems with filter bank position 9B and 4A. Bad relays were found and replaced.Filter position 4A was fixed but filter position 9B still with problems.
crvTuesday, April 15, 2008 11:34:55Filter 9 problem was fixed. We replaced the four RF switches for two new switches.
crvTuesday, April 15, 2008 11:37:07Problem with cal values wer fixed thigthning postamp chasis internal cabling. LBW system was recalibrated yesterday Monday april de 14th 2008
crvWednesday, April 30, 2008 10:56:46Dewar was warmed during the weekend of april 27th. The dewar was opened on monday april 28th and the and the amplifiers,couplers and charcole trap bought down for servicing.
crv.Wednesday, April 30, 2008 11:01:05Amplifiers were checked and CH B amp. L128 was replaced for the spare L127. The traps charcole was replaced for new activated charcole because the original epoxy was loose. The amps were installed and the dewar was closed left doing vacuum. april 29th
crvMonday, June 23, 2008 08:32:40lbw crosshead ed 29 mechanicaly failed and the system was warmed up on friday june the 20th 2008. On sunday June the 22th the crosshead was replaced and started cooling.
crvThursday, July 17, 2008 16:52:53On Monday july the 7th 2008 the system was shut off because the cryo pump was having mechanical problem. The system will be cheked and cleaned .
crvThursday, July 17, 2008 16:55:54Monday july 12th 2008 LBW compressor and cryo lines were connected to the test dewar to trap any contamination in the system. Thursday july 17th 2008 the dewar didn,t show any contamination problem and was disconnected.
crvThursday, July 17, 2008 16:58:26CBH cryo pump was removed cleaned and installed on LBW to cooldown the reciever. Temperature at the time of leaving the dome were: 16k =19K 70k= 137.2k omt= 274.4
crvFriday, July 18, 2008 14:32:10dewar temps are 16k=13.5 70k= 66.5 omt= 100.3 . Cryo pump is sounding good no racheting or sing of contamination.
crvMonday, August 4, 2008 16:35:47Lbw rfi problem not detected with the spectrum analizer still the cryo pump driver was removed shielded tested and installed back on LBW. System to be checked by Barbara Catinella.
Dana Whitlowdwhitlow@naic.eduFriday, October 3, 2008 09:18:5131 Aug 08: Took LBW out of service to address intermittent RFI spikes which were plaguing observations. Opened up dewar. Next few days: Found flaky connections at inside of vacuum-tight connector for lines to bias chassis. Replaced cable between the connector and the barrier-type terminal strips inside dewar. Also found could generate noise spikes by striking wiring to dewar heater, even with dewar heater circuit completely isolated from system. Suspect that this was due to tiny ESDs on plastic insulation on wiring, and decided that this was not relevant to normal system operation. 5 Sept 08: Buttoned up system and re-cooled dewar. 7 Sept 08: Dewar (incl OMT) fully cooled.
Dana Whitlowdwhitlow@naic.eduFriday, October 10, 2008 08:53:41New CAL values were installed 26 Sept 08. However, the accuracy of these has since come into doubt and they will be re-measured soon.
crvTuesday, October 21, 2008 16:57:09reciever cryo pump was sounding contaminated. The reciever was warmed up to 125K and the cryo pump was purged. Reciever started cooling down at 15:00 .
Dana Whitlowdwhitlow@naic.eduMonday, December 15, 2008 14:25:36The suspicious CAL values from 26 Sept 2008 were replaced by newly-determined values on 15 Dec 2008. These look much better.
Stephen Jensensjensen@naic.eduWednesday, February 18, 2009 12:51:37Had to do a 100K purge for LBW as it was starting to ratchet.
crvcrios@naic.eduThursday, May 7, 2009 16:36:54The reciever is warming up to replace the cryo pump. The cryo pump will be changed to see if the RFI that is affecting barbara Catinella stops. The cryo pump will be replace tomorrow May 8th 2009.
crvFriday, May 8, 2009 14:16:56The cryo pump was replaced for a new one and the dewar started cooldown at 11:45 AM.
crvThursday, August 6, 2009 15:15:18reciever was vacuumed and is cooled down today 8/3/2009 after sunday night power failure and wednesday power failure .
criosFriday, April 30, 2010 17:37:10A switch was connected to the Amp. LED circuit for troubleshooting purposes.
crvcrios@naic.eduThursday, August 26, 2010 15:05:27on august 19th 2010 Lbw blew away the vacuum window from the dewars OMT and had to be brought down for service. Dewar lowered from dome on August 20th 2010.
crvMonday, September 27, 2010 13:47:55system was installed and cooling on friday september 27 2010. System is pending test.
crvcrios@naic.edu.Monday, September 27, 2010 16:59:54cal ware not switching and generated bad tsys values. Chassis of CBH were not powered but connected to the mux chassis . Chassis were powered and the cal started to switch correctly. Tsys values normal.
crvcrios@naic.eduThursday, October 28, 2010 13:06:00The system cryogenics where turned off on Tuesday Oct. 26th 2010 due to cryo genics maintenance.
crv.crios@naic.eduThursday, October 28, 2010 13:12:12The system started cool down at 10:30am. At the moment of starting the cooldown the vacuum was on 3 microns. The displacer and cycinder were clean no aparent contamination. We readjusted the displacer clearence and the nocking sound is gone.
crvcrios@naic.eduFriday, October 29, 2010 11:53:44The vacuum valve on the dewar was closed today and the receiver is ready to be used.
criosCrios@naic.eduTuesday, December 7, 2010 10:28:22The system is warming up since yesterday night. The cryo pump will be checked do to mechanical noise. and the dewar will be vac.do to last week power failure.
crvcrios@naic.eduWednesday, December 8, 2010 13:47:47The cryo pump was cleaned and purged. The dewar vac. was o.k. and the system started cooldown.
crvcrios@naic.eduTuesday, December 28, 2010 11:36:29Lbw was turned off by Ganesh and Dana on sunday Dic 26th. on monday Dic. 27th the cryo pump was replaced by Carlos and Alfredo. Today We started the cooldown after letting it vac.pump for a day. We purged the the complete system compressors, lines and cryo pump before the cooldown.
crvcrios@naic.eduWednesday, December 29, 2010 10:47:30the receiver is cold and the cryo pump is running smooth. vac. pump removed.
crvcrios@naic.eduTuesday, February 1, 2011 09:09:00While checking for the stability problem on ch B of lbw we found that the Type N connector on the 1320 filter had a bad solder joint. The connecter could be removed from the cable the center pin was fine . We resolder the connector. system pending to be tested.
Dana Whitlowdwhitlow@naic.eduTuesday, April 19, 2011 17:53:04LBW started cooling approx 17:10 19 April 2011, following repairs to crosshead arising from yesterday's failure.
Dana Whitlowdwhitlow@naic.eduThursday, April 21, 2011 08:10:53On Tuesday May 19 a leak rate check was performed on LBW shortly before re-cooling was started. The rate was 7 microns over 10 minutes. This extrapolates to 1 atm over about 2 years.
Dana Whitlowdwhitlow@naic.eduThursday, April 21, 2011 08:26:02CORRECTION TO PREVIOUS ENTRY: The leak rate check was performed on April 19, not May 19.
crioscrios@naic.eduWednesday, May 4, 2011 14:00:00post amp and polarization box bypass chassis was installed on the LBw . This new chasis will bypass the polarization box and let you use a 1550MHz low pass filter. The chasis was constructed using the CBH polarization box and will be controlled with the CBL polarization box control bit. The change was made for a Barbara catinella experiment.
Carlos Rioscrios@naic.eduThursday, February 2, 2012 10:20:28The dewar safety pressure plug poped out and the system lost it's vacuum.System was turned of and started warmup. The dewar is doing vacuum. Feb. 1st 2012.
Dana Whitlowdwhitlow@naic.eduFriday, May 4, 2012 14:07:16BRIEF REPORT ON AO RADAR BLANKER PROBLEM SEEN 03/04MAY2012 LBW observer complained that the radar blanker was not doing its job, and it seemed not to be locking. First obvious symptoms noted: the unlock light was flashing at a regular rate (~2 Hz?), the VCO tuning voltage was indicating at or near the high end of its range, and the "acquired signal level" was reading near zero. A quick check with a 'scope revealed that the IF signal looked very healthy and that the 5-wide radar pulse sequence looked normal, but the blanking pulse train was a bit late and was sort of oscillating over time positions ranging from a tiny bit late to several hundred microseconds late. The oscillation period was somewhat variable, but was typically a modest fraction of one second. I captured ~10 sec of data from the "detected output" port of the radar blanker with a DPO4104 with 10,000,000 record length and did a Fourier analysis, and found a sequence repetition rate of 355.72 Hz. But then, unbeknownst to me, Phil called the radar site to inquire, and checking revealed a COHO lock loss at the radar. They relocked it, and our radar blanker began working properly again. I redid the 'scope measurement and now found a sequence repetition rate of 354.53 Hz, which is about 0.36% lower. So next time this happens, check the SRR and if it's not well within 0.3% of 354.528Hz (probably should be well within 0.1%), blame the radar first. Dana Whitlow 04 May 2012
Denis Urbaindurbain@naic.eduWednesday, October 31, 2012 17:27:22The Pol A RF signal was unstable, Phil found that the problem was present only when the High Pass Filter 1370 was used. The two HP 8762A Coaxial switch that connect to the filter have been replaced by Wilson Morales.
Denis Urbaindurbain@naic.eduTuesday, November 20, 2012 13:32:53The Helium cooling circuit has developed a leak. Over the weekend the pressure drop reduced the cooling capacity of the refrigerator. The loss of cooling power increased the inner temperatures, the increase in temperature worsen the vacuum and started a runaway cycle that ended with the overpressure cap being pushed out and total loss of vacuum. The compressor was stopped, the overpressure plug was re-inserted in position and 2 vacuum pumps were connected to restore the vacuum after the blowout. After 8 hours of pumping and to meet the observation schedule the cooler was started. The leak was found in one of the cable wrap line and the line was removed.
Denis Urbaindurbain@naic.eduFriday, December 7, 2012 16:44:51The L band wide receiver was brought down to the lab to investigate the high Tsys and the 16K stage temperature readout problem. The receiver was opened but no trace of water contamination was found inside. The OMT and the small shield around the LNA were wrapped with super-insulation. One of the RF cable (pol A) had a small crack at the solder joint of the SMA connector. Both RF cables were touched up. The Charcoal trap was removed because the epoxy was braking off and pieces of charcoal were falling off. Guarionex repaired the charcoal trap but it could not be installed because the epoxy used was not dry on time. The complete receiver was clean and reassembled. The temperature reading problem was troubleshooted by Carlos and Wilson and the 5V power supply from the post amp chassis was replaced. The LNA’s were measured warm and the gain was as expected but the noise temperature was higher than the NRAO specification. We tried another amplifier that gave a similar temperature therefore we decided to leave the LNA untouched. The receiver was closed and pumped down, the coldhead was cleaned and the displacers baked overnight in the vacuum oven. The vacuum was around 1 10E-2 Mbar in the morning, the superinsulation might be the reason for the rather high pressure. Leak testing the receiver didn’t show any problem. The pressure is going down but slowly.
Denis Urbaindurbain@naic.eduMonday, February 25, 2013 14:07:49The colhead was making a ratcheting noise and was turned off for service. Water was found on the displacer when the coldhead was removed. The vacuum was pretty bad and two vacuum pumps were connected in parallel with a T. After about 24h, the pressure dropped around 20 millitorr which is quite high. The RF window was inspected but it looked fine. The Dewar should be leak tested in the lab to understand the reason for the fairly high vacuum pressure.
Denis Urbaindurbain@naic.eduThursday, June 6, 2013 08:25:59The coldhead was contaminated, the receiver was warmed up on the 5/30/2013. The next day the cold head was cleaned, the vacuum was restored (2 vacuum pumps were connected). The same day the receiver was cooled again because of the observinf schedule.
Anthony Fordaford@naic.eduThursday, October 31, 2013 16:40:59Tsys measurements were showing anomalous readings on certain cal modes (noticed on 10/30/13, backtracking reveals possible anomalies as far back as 10/22/13). After testing logic board and HP8762A relays (RL 3 & 4), it was determined they were not at fault. Further investigation showed that the Vcc on the enable relay after noise diode A was shorting to the case. The connections were re-soldered to prevent shorting and the problem was resolved.
Felix O. Fernandezfelix.fernandez@naic.eduTuesday, December 24, 2013 14:44:23The LBW Rx's cold-head has been "knocking" for the past few weeks, thus time was allocated to service the cold-head. On DEC-22-2013 the LBW Rx was warmed up. The replacement cold-head was installed on DEC-23-2013. Finally, the cool down process was started on DEC-24-2013.
Cold-head replacementfelix.fernandez@naic.eduTuesday, March 18, 2014 08:03:59Since MAR-11-2014 the LB OMT has been warming up. On MAR-12-2014 the OMT had reach a temperature of 180K, well beyond the operational level. The LBW was warmed-up on MAR-12-2014. Here's the progress. -MAR-12-2014: The pop-off valve was expelled during the warm-up process. -MAR-13-2014: Replaced cold-heads and started vacuum process using the large scroll pump and a typical large turbo pump. -MAR-14-2014: Vacuum process had stalled at 24 mTorr. After further inspection it was found that the pop-off valve had a vacuum leak. After re-installation of the pop-off valve the vacuum process proceeded beyond 20 mTorr. -MAR-15-2014: The vacuum had stalled at 10mTorr. It seem that the large scroll pump has a check valve that levels off at 10mTorr. Once the large pump was removed the vacuum process continued below 10 mTorr. -MAR-17-2014: The cool-down process started with a vacuum level of 6.5 mTorr.
Cold-head decontaminationfelix.fernandez@naic.eduTuesday, March 18, 2014 08:07:44On MAR-17-2014 during the morning hours, about 09:00, the LBW cold-head started knocking due to contamination. The LBW started the warm-up process and at 18:00 the cold-head was purged and re-installed and the cool-down process started once again.
Cold-head contamination or mechanical glitchfelix.fernandez@naic.eduTuesday, March 18, 2014 15:08:08During today's inspection of the cool-down process (which started yesterday evening) the cold-head was found to be "knocking". Since strict procedures were followed during the cold-head's installation it seems very unlikely that the cold-head got contaminated and, therefore, the culprit of this situation seems to be of mechanical nature and not contamination. The LBW has started the warm-up process at 09:00 and it is scheduled for cold-head replacement tomorrow (MAR-19-2014) morning.
Carlos Rioscrios@naic.eduThursday, March 20, 2014 15:19:56Cooldown started yesterday at 10:00am. The system is cold and ready to be used.
Vacuum leak?felix.fernandez@naic.eduTuesday, August 26, 2014 15:32:37LBW has been increasing in temp for the last few weeks. After a vacuum check the gauge read 2.5 mTorr where the typical values should be in the order of 1e-5 (i.e. readings below zero when using our Teledyne Hastings DV-5 thermocouple vacuum gauge tubes). This decrease in vacuum increases the thermal load due to convection within the receiver. We connected a vacuum pump yesterday, trying to re-achieve proper vacuum. After 24Hrs the vacuum readings where within specs and the receiver's OMT temperature went from 130K to 81K. Conclusion: LBW might be suffering from a small vacuum leak. Hence, the receiver's vacuum must be monitored on a regular basis.
Contamination -Thermal Cycle-Compressor Exchangefelix.fernandez@naic.eduFriday, September 26, 2014 11:35:10LBW cold-head was ratcheting on Tuesday SEP-23-2014. It was decided to warm up the system for servicing. The contamination was later found to be humidity inside the cold-head. The cold-head displacers were baked and clean. The dewar is now at 2mTorr and ready for cool-down. We expect to start the cool-down tomorrow during the Dual-Beam setup. Furthermore, the compressor went dead on Wednesday SEP-24-2014. The compressor was replaced with out M450 spare compressor.
Ch B tsys values high.crios@naic.eduMonday, October 6, 2014 14:39:39Tsys on ch B was high. The cable from dewar rf outputB to post amp inputB was retighten. Tsys values are normal.
LBW Cryo Monitor cable replacedfelix.fernandez@naic.eduTuesday, October 21, 2014 15:07:45LBW has been having issues with its cryo monitoring system. The problem is that LBW's temperatures oscillate drastically while Tsys values show a perfectly healthy system. Today after checking several possible candidates the Cryo Monitor cable (thin flexible sliver cable) was found to have its shield separated from the connector. We replaced the cable with the spare one in hopes to correct the cryo monitoring issues.
Cryo Monitor issues persistfelix.fernandez@naic.eduMonday, November 10, 2014 10:59:51LBW's temperature readings keep oscillating over the past few days. Today we found out that if you manipulate the ANALOG MONITOR (P2) cable the temperature reading change dramatically. The cable was inspected and deemed to be OK. The only other option is the post-amp chassis connector. We could not check this chassis today due to SB Klystron repairs. We will check tomorrow for problems inside the post amp chassis.
ANALOG MUX card replacementsfelix.fernandez@naic.eduWednesday, November 12, 2014 11:33:57LBW's ANALOG MUX (AO940330) was replaced today in an effort to suppress the temperature reading oscillations. After the replacement the temperatures seem stable but not accurate. The 16K stage readings are about 8K and the OMT are about 23K. Although these readings are off they are stable enough for us to evaluate the cryogenic system's status.
Relay #4 replacement (Tsys readings related)felix.fernandez@naic.eduWednesday, November 12, 2014 11:37:54There has been some unusual readings been reported for corcal and xcal on both the high and low setting for Pol-B. The system values differ by more than 20K, this is not typical. We decided to replace relay #4 (HP8762A) which controls the signal path to Pol-B. After the replacement the Tsys readings are within 2K a typical situation. We will evaluate the relay with the network analyzer in the near future.
OMT temp increasefelix.fernandez@naic.eduTuesday, December 9, 2014 14:48:44LBW's OMT temperature had increased to 140K overnight, which is much higher than the typical 70-80K. We found that LBW has lost some of its vacuum and was at a 3 mTorr vacuum. A vacuum pump was added to reverse and mitigate the temperature increase. We think that most likely the compressor went "dead" for a couple of minutes, maybe due to overheating problems. We will keep LBW under close observations.
Temperature increase and vacuum resetfelix.fernandez@naic.eduWednesday, December 24, 2014 10:58:02Once again, LBW's temperature increased and vacuum became marginal. Once vacuum was reset (DEC-23-2014), the system temperature became nominal (DEC-24-2014). NOTE: We need to service LBW soon.
LBW maintenance 19-JAN to 25-JANfelix.fernandez@naic.eduMonday, January 26, 2015 13:54:48During the last DB setup we had the opportunity to service LBW and try to repair its vacuum leak. The actual leak was found at the receiver's pop-up valve. A new pop-up valve was fabricated to include a thicker o-ring. After the new pop-up valve was installed the system was evacuated and cooled down. The systems seem to be behaving without any trace of a vacuum leak, but only time will tell. Also, during this maintenance window the current vacuum gauge on the Dewar was deemed faulty and replacement is needed.
Thermocouple Vacuum Gauge Tube Replacementfelix.fernandez@naic.eduTuesday, February 10, 2015 11:44:07On Feb-09 LBW was warmed up to replace the vacuum gauge. Today, the gauge was replaced and the evacuation process was started.
Polm. B jumpcrios@naic.eduThursday, April 9, 2015 09:55:27We found that the pol B output cable of the dewar, and the pol B cable from the lbw filter bank to the blue cabinets were causing problems. After tightening the cables, the problems stopped.
Cold-head replacementfelix.fernandez@naic.eduFriday, May 8, 2015 09:43:33On MAY-6 we replaced the cold-head and found humidity inside the old cold-head. A "fresh" cold-head was installed and cool down started on MAY-8 with a starting vacuum of 2.5 mTorr. LBW should be ready by MAY-9
PolB LNA and Vacuum windowfelix.fernandez@naic.eduFriday, May 15, 2015 10:45:38During the last warm-up cycle as LBW was being evacuated stage 2 of pol-B's LNA became unstable and its drain current rose to 25mA (MAY-7 10PM). The LNA never recuperated from this event. Furthermore, the next Monday (MAY-11) as we were disassembling LBW's feed-horn we found out that the vacuum window was close to a complete failure. The vacuum window's foam had dislodged from its position and the Kapton film become convex. Basically LBW's vacuum window almost imploded. The old LNA (L127) has since been replaced by another LNA from NRAO (L004) and the vacuum window was sent for repairs to NRAO, NM.
LBW returns to the Domefelix.fernandez@naic.eduFriday, June 5, 2015 11:51:55LBW was brought up to the Dome today. The vacuum was repaired by NRAO, NM(twice, since the first repair also failed during evacuation) and replaced. The foam was replaced with the same material as EVLA's and added one extra layer of foam for IR filtering just like the EVLA L-band receiver. We will start the cool-down today. Also, the cable from the input of both 1320 HPF were damaged (broken shield at N-connectors) so they were replaced by SMA-SMA cables and SMA 90-elbows.
Cryopump with racheting noise.crios@naic.eduThursday, July 9, 2015 12:54:49The cooling system of LBW was turned off due to a racheting noise. The system will be checked after the 430MHz world day runs.
lbw He line and cryopump changecrios@naic.eduTuesday, September 22, 2015 14:00:42Cryo lines were change on friday sept. 11 to clean the original lines from any posible oil contamination. This was made to solve the varios cryopump change made to LBW in the las couple of weeks. A new cryo pump was installed and started cooling
Cryo pump torque problem.crios@naic.eduTuesday, September 22, 2015 14:10:09On sunday sept. 13 the cryopump was turned off do to a sound similar to conrtamination. All the components in the system are new and free of contamination. After further troubleshooting the problem is that the cryo pumps have poor torque and other receiver are also affected. The cryo pump power supply output are of 160V ac and the mininal working voltage should be 180V ac. The input voltage for the power supply is 190V ac and not 208V ac and compressors voltage is 450V ac not 480V ac. waiting for voltage issue to be fix on the week.
Cryo pump timing check.crios@naic.eduTuesday, September 22, 2015 14:14:29Do to the low torque issue we re ajusted the timing of the valves. the pump was timed to have less time the input valve open. we expect that having less volume of He in the cooling chamber will help the low torque problem. Cryo pump installed and cooling sounds fine.
Dome voltage work. Tuesday, September 22, 2015 14:22:35On Wednesday Sept. 15 The LBW cryopump sounding fine no low torque sound. Compressor running at 245 psi minimal efective pressure for cooling the dewars. Maintenance crew working on the Dome low voltage problem. They adjusted the tap on the main transformer twice at the end of the day we had 208V ac on the dome and 460V ac on the compressor. Compressor voltage still to be solved . Cryopump power supply output voltage at 180V ac. All the cryopump motor sound improve.
Pol B. drifting Check.crios@naic.eduTuesday, September 22, 2015 14:25:27At this moment the Receiver is working great. Phil made a test by turning the cryo off and he concluded that Ch B drift as it warms proble shoul be in the dewar.
Compressor Failurefelix.fernandez@naic.eduTuesday, June 7, 2016 14:32:40During last weeks HF campaign the LBW's compressor failed The comprssor shut-down and tripped the circuit brakes in the distribution panel. After and internal inspection it was found that the contactor input terminal had "melted" due to a flash event. After replacing the damaged contactor and overload relay (now set at 13Amp) the compressor is now back in working order.
Replaced RL3 RF switch felix.fernandez@naic.eduFriday, June 2, 2017 13:09:14During the past weeks LBW Tsys has been showing higher than usual readings on PolA for the following cals: hxcal, h90cal, lxcal, l90cal. Upon inspection it was decided to replace RL3 in the post-amp chassis. After this the Tsys values are back to normal.
Cool-down started`felix.fernandez@naic.eduThursday, October 26, 2017 13:00:01Cool-down started today at 9:00AM. LBW should be ready within the next 36hrs.
Cold-head replacementfelix.fernandez@naic.eduFriday, January 19, 2018 11:08:49LBW cold-head was replaced during Jan 15-19 430Tx campaing. This was in part to recent reports of gain variation on Pol-A during several VLBI runs. The problem was deemed to be before the system's hybrid used for circular polarization observations. Furthermore, the cold-head was due for service after almost three years of use. The whole thermal cycle was completed today.
Pol-B LNA Deadfelix.fernandez@naic.eduTuesday, January 30, 2018 14:26:51LBW pol-B is dead. Today we went to the Dome and plugged the spectrum analyzer to the dewar's output. Pol-B showed a flat line while Pol-A had some pass band gain. LNA needs to be replaced.
LBW back to normal statefelix.fernandez@naic.eduTuesday, February 6, 2018 08:15:21LBW was thermal cycled (unintentionally) during last week's "dual-beam" observation. This thermal cycle "fixed" the problem with Pol-B. Hence, the problem seems to be a intermittent physical connection within the dewar. This problem is expected to return at some point.
LBW warmed upfelix.fernandez@naic.eduWednesday, January 16, 2019 10:20:51On 12-jan-19 there was a power dip and LBW started an unexpected warm-up. After this warm-up Pol-B's LNA stage #3 current started to increase. It seems the gate of stage 3 has become "loose". We need to warm-up and open LBW to investigate.
PolB 1370 HPF felix.fernandez@naic.eduTuesday, April 23, 2019 13:26:46The 1370 HPF on pol-b was stuck in an "always on" state. We brought down the filter bank chassis and found no problems. We then replaced both HP8762A switches that enable/disable the 1370 HPF but the problem remained. The final "solution" was to swap the control cable from pol-a to pol-b. After the cable swap the system works as expected. NOTE: if you try (if possible) to change the 1370 HPF state on JUST pol-a, you will now change the state on pol-b, since the control cables are now swapped.

<> Back to Top